# arXiv:2402.18042v3[math.AG] 01 Mar 2026

### v3

@@ 1 Introduction
Theorem 1.1.
(cf. Theorem 2.26) Let $A$ be the henselization of the local ring $\mathcal{O}_{Y,y}$ of a normal crossing variety $Y$ at a point $y\in Y$. Then the sequence
@@
  • Property $\operatorname{P}_{3}(a)$: Let $A$ be the henselization of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point of $y\in Y$. Then the sequence
@@
By Remark 2.12, $(\mathcal{F}_{X},N)=(\lambda^{n}_{X,r},1)$ satisfies Condition 2.11 and so Theorem 1.1 holds for $u\geq 1$ by Property $\operatorname{P}_{3}(a)$ of Theorem 1.2. Then we are able to prove Theorem 1.1 for $u=0$ by using the coniveau spectral sequence (1). This concludes the outline of the proof of Theorem 1.1.
@@
(cf. Corollary 2.23) Let $A$ be the henselization of the local ring $\mathcal{O}_{Y,y}$ of a normal crossing variety $Y$ over the spectrum of a field $k$ at a point $y\in Y$ and $l$ an integer which is prime to $\operatorname{char}(k)$. Then the sequence
@@
In the proof of Theorem 1.2, we use Lemma 2.21. The proof of Theorem 1.2 is similar to that of Lemma 2.21 and we prove them by downward induction on $N^{\prime}$ and induction on the dimension of a scheme. We use Propositions 2.7 and 2.9 to apply downward induction on $N^{\prime}$ and we are able to prove them by using Lemma 2.2. We use Lemma 2.19 to apply induction on the dimension of a scheme. See §2.1 below for more detailed outline of the proof.
@@
    Let $i:Z\to\mathfrak{X}$ be a closed immersion of codimension $1$ and $Z$ an irreducible closed scheme of $\mathfrak{X}$ with $\operatorname{char}(Z)=p$. Then the canonical map induces a quasi-isomorphism
@@
For a scheme $X$, $X^{(s)}$ denotes the set of points of codimension $s$ and $k(X)$ denotes the ring of rational functions on $X$. For a ring $A$, $k(A)$ denotes $k(\operatorname{Spec}(A))$. For a point $x\in X$, $\kappa(x)$ denotes the residue field of $x$, $\mathcal{O}_{X,x}$ denotes the local ring of $X$ at $x\in X$ and $\mathcal{O}^{h}_{X,x}$ denotes the henselization of $\mathcal{O}_{X,x}$.
@@ 2 Gersten-type conjecture
for some integer $a$ with $0\leq a\leq M=\operatorname{dim}(Y)$ (cf. [28, pp.180–181, Definition 2.1]). In this section, we prove the Gersten-type conjecture for a pair of a sheaf $\mathcal{F}$ on $Y_{\mathrm{\acute{e}t}}$ and a non-negative integer $N$ which satisfies Condition 2.11 over the henselization ring of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point $y\in Y$ ( Theorem 1.2, Theorem 2.22). Let
@@
The outline of the proof of Theorem 1.2 is as follows: By [17, p.93, III, Corollary 1.28], it suffices to prove $\operatorname{P}_{1}(a)$ and $\operatorname{P}_{2}(a)$ in the case where $Y$ is replaced by the spectrum of the henselization $A=\mathcal{O}^{h}_{Y,y}$ of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point $y\in Y$ and a point $y\in Y$ is replaced by the maximal ideal of $A$. We put $\operatorname{Spec}(C)=\operatorname{Spec}(A)\times_{Y}Z$ and prove Theorem 1.2 by induction on $a=\#(\operatorname{Spec}(A)^{(0)})$.
@@ 2.2 Equi-characteristic cases
In this section, we frequently assume that a ring $A$ is a formally equidimensional noetherian local ring (cf. [16, p.251, Definition]). For example, an equidimensional local ring which is essentially of finite type over a regular local ring is formally equidimensional by [16, p.251, Theorem 31.6 (iii)].
@@
Let $A$ be a formally equidimensional noetherian local ring. By [16, p.251, Theorem 31.6 (iv)] and [16, p.250, Theorem 31.5], AA is equidimensional. By [16, p.251, Theorem 31.6 (i)], $A_{\mathfrak{p}}$ is also formally equidimensional for any $\mathfrak{p}\in\operatorname{Spec}(A)$. Moreover, the henselization $A_{\mathfrak{p}}^{h}$ of $A_{\mathfrak{p}}$ is formally equidimensional by the definition.
@@
Let $A$ be a formally equidimensional noetherian local ring, $\mathcal{F}$ a sheaf of groups on $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$ and $x\in\operatorname{Spec}(A)^{(s)}$. Put
@@
By Remark 2.3 and [17, p.93. III, Corollary 1.28], it suffices to prove the statement in the case where $A$ is a formally equidimensional henselian local ring and $x$ is the maximal ideal of $\operatorname{Spec}(A)$. We prove the statement by induction on $\operatorname{dim}(A)$.
@@
Let the notations $A$, $\mathcal{F}$ and $T$ be the same as in Proposition 2.4. Suppose that $\operatorname{dim}(A)=s$. Then $\operatorname{H}^{t}_{\mathrm{\acute{e}t}}(A,\mathcal{F})$ is a torsion group for $t\geq s+T$.
@@
Let $A$ be a henselian local ring, $I$ an ideal of $A$ and $\mathcal{F}$ a sheaf of groups on $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$. Then we have an isomorphism
@@
Let $A$ be a formally equidimentional noetherian local ring over a field $k$, $l$ a prime number which is prime to $\operatorname{char}(k)$ and $\mathcal{F}$ a sheaf of groups on $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$. Let the notation $T$ be the same as in Proposition 2.4. Suppose that $\operatorname{cd}_{l}(k)<\infty$ and $\operatorname{tr.deg}_{k}A_{x}<\infty$ for any $x\in\operatorname{Spec}(A)^{(0)}$. Put
@@
It suffices to prove the statement in the case where $A$ is a formally equidimensional henselian local ring and $x$ is the maximal ideal of $\operatorname{Spec}(A)$. We prove the statement by induction on $\operatorname{dim}(A)$.
@@
for $v\geq N+T+1$. In order to prove the statement, we use Lemma 2.2. Since $A$ is henselian, we have an isomorphism
@@
for $m\geq 0$ where $i:\operatorname{Spec}(\kappa(x))\to\operatorname{Spec}(A)$ is the closed immersion. Since $A$ is universally catenary by [16, p.251, Theorem 31.6 (iv)], we have
@@
Let $A$ be a formally equidimensional local ring of positive characteristic $p>0$ and $\mathcal{F}$ a sheaf of groups on $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$. Since $\operatorname{cd}_{p}(A)=1$, we have
@@
Let $A$ be a formally equidimensional local ring and $\mathcal{F}$ a sheaf on the étale site $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$. Let $N$ be a non-negative integer and $l$ a prime number. Suppose that $\operatorname{dim}(A)=s>1$ and $N\geq T$ where the notation $T$ is the same as in Proposition 2.4.
@@
for $s-1\leq e\leq s$. We prove the isomorphism (18) by using Lemma 2.2.
@@
Let $\mathcal{V}$ be the category whose objects are normal crossing varieties over the spectrum of a field $k$ and whose morphisms are morphisms of schemes. Let $X$ be an object in $\mathcal{V}$, $\mathcal{F}_{X}$ an étale sheaf of abelian groups on $X$ and $f:X\to Y$ a morphism in $\mathcal{V}$. Let $N$ be a non-negative integer. Then $(\mathcal{F}_{X},N)\in\mathbb{S}_{X_{\mathrm{\acute{e}t}}}\times\mathbb{N}_{\geq 0}$ satisfies the followings:
@@
Let $X$ be essentially of finite type over the spectrum of a regular ring of dimension at most $1$, $n$ an integer and $\mathbb{Z}(n)$ Bloch’s cycle complex for the Zariski topology. Let $i:Z\to X$ be a closed immersion of codimension $c$ with open complement $j:U\to X$, $i_{1}:W\to X$ a closed immersion of codimension $c^{\prime}$ and
@@
is exact for any integer $i$. If $\operatorname{char}(k)=p\geq 0$, then we have a quasi-isomorphism
@@
for any integer $i$. So the isomorphism (27) follows from Lemma 2.1. This completes the proof. ∎
@@
Let $Y$ be a smooth scheme over the spectrum of a field $k$ and $Z$ a normal crossing variety over $\operatorname{Spec}(k)$ which is a closed scheme of $Y$ of codimension $1$.
Let $A$ be the henselization of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point $y$ of $Y$, $C$ the henselization of the local ring $\mathcal{O}_{Z,y}$ and $i:\operatorname{Spec}(C)\to\operatorname{Spec}(A)$ the corresponding closed immersion.
@@
We prove the injectivity of the homomorphism (37) for $t=N^{\prime}$ by induction on $s$. By [17, p.93, III, Corollary 1.28], it suffices to prove the injectivity of the homomorphism (37) in the case where $A$ is a henselian regular local ring with $\operatorname{dim}(A)=s$.
@@
In the case where $s=1$, we prove the injectivity of the homomorphism (37). Since $A$ is a henselian valuation ring, we have an isomorphism
@@
  • Property $\operatorname{P}_{3}(a)$: Let $A$ be the henselization of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point of $y\in Y$. Then the sequence
@@
First, we prove $\operatorname{P}_{1}(a+1)$. By [17, p.93, III, Corollary 1.28], it suffices to prove $\operatorname{P}_{1}(a+1)$ in the case where $Y$ is the spectrum of a henselian local ring $A$ and $\operatorname{dim}(Y)=s$. Since $(Z^{(0)})=a$, $\operatorname{P}_{1}(a+1)$ follows from $\operatorname{P}_{3}(a)$ and Lemma 2.18.
@@
is injective in the case where $A$ is the henselization at a point $y$ of $Y^{(s)}\cap Z$ by [17, p.93, III, Corollary 1.28]. Since we have
@@
Let $A_{a+1}$ be the henselization of $\mathcal{O}_{Y_{a+1},y}$ and $U^{\prime}=\operatorname{Spec}(A_{a+1})\times_{Y_{a+1}}U$. For notational simplicity, a base change of $j$ (resp. $j^{\prime}$, $i$) is also denoted by $j$ (resp. $j^{\prime}$, $i$), unless confusion arises. Then we have an isomorphism
@@
Finally, we prove $\operatorname{P}_{3}(a+1)$. Let the notations $A$, $C$ and $U^{\prime}$ be the same as above. By (2.2) and $\operatorname{P}_{3}(a)$, the sequences
@@
are exact for $N^{\prime}\geq N_{1}$. Since $A$ is henselian, we have isomorphisms
@@
for $u+v=s^{\prime}+t$. Hence, the statement follows from Lemma 2.2. This completes the proof.
@@
Let $A$ be the strict henselian of a local ring of a normal crossing variety over the spectrum of a field of positive characteristic. Then $A$ is a direct limit of a family of the strict henselian rings $A_{i}$ ($i\in I$) of a local ring of a normal crossing variety over the spectrum of the prime field which satisfies the following:
Let $A/(a_{s})$ $(s\in S)$ be the irreducible components of $A$. Then
@@
where $k_{i}$ run over subfields of $k$ which contain $k^{\prime}$ and are finitely generated over the prime field. Let $\mathfrak{m}$ be the maximal ideal of $C$, $f_{i}:k_{i}\otimes_{k^{\prime}}C^{\prime}\to C$ the base change and $\mathfrak{m}_{i}=(f_{i})^{-1}(\mathfrak{m}).$ Put $C_{i}=\left(k_{i}\otimes_{k^{\prime}}C^{\prime}\right)_{\mathfrak{m}_{i}}.$ Then $C_{i}$ is regular by [16, p.182, Theorem 23.7 (i)]. Let $c^{\prime}_{s}$ ($s\in S$) be elements of $C$ such that $C/(c^{\prime}_{s})\simeq A^{\prime}/(a^{\prime}_{s})$ where $A^{\prime}/(a^{\prime}_{s})$ ($s\in S$) are the irreducible components of $A^{\prime}$. Since we are able to choose a directed subset $J$ of $I$ such that $c_{s}^{\prime}\in\mathfrak{m}_{j}$ and
@@
for any $i\in I$ and $A$ (resp. $A_{i}$) denotes the strict henselian of $A^{\prime}$ (resp. $A^{\prime}_{i}$). Then $A$ is also the strict henselian of $\mathcal{O}_{Y,y}$ and $A_{i}$ satisfies the property of the statement. This completes the proof. ∎
@@
Let $A$ be the henselization of the local ring $\mathcal{O}_{Y,y}$ of a normal crossing variety $Y$ over the spectrum of a field of positive characteristic $p>0$ at a point $y\in Y$. Then the sequence

### v4

# Gersten-type conjecture for henselian local rings of normal crossing varieties

@@ 1 Introduction
Theorem $1.$
(cf. Theorem 2.26) Let AA be the henselization of the local ring $\mathcal{O}_{Y,y}$ of a normal crossing variety $Y$ at a point $y\in Y$. Then the sequence
@@
  • Property $\operatorname{P}_{3}(a)$: Let AA be the henselization of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point of $y\in Y$. Then the sequence
@@
By Remark 2.12, $(\mathcal{F}_{X},N)=(\lambda^{n}_{X,r},1)$ satisfies Condition 2.11 and so Theorem 1.1 holds for $u\geq 1$ by Property $\operatorname{P}_{3}(a)$ of Theorem 1.2. Then we are able to prove Theorem 1.1 for $u=0$ by using the coniveau spectral sequence (1). This concludes the outline of the proof of Theorem $1.$
@@
(cf. Corollary 2.23) Let AA be the henselization of the local ring $\mathcal{O}_{Y,y}$ of a normal crossing variety $Y$ over the spectrum of a field $k$ at a point $y\in Y$ and $l$ an integer which is prime to $\operatorname{char}(k)$. Then the sequence
@@
In the proof of Theorem 1.2, we use Lemma 2.21. The proof of Theorem 1.2 is similar to that of Lemma 2.21 and we prove them by downward induction on $N^{\prime}$ and induction on the dimension of a scheme. We use Propositions 2.7 and 2.9 to apply downward induction on $N^{\prime}$ and we are able to prove them by using Lemma $2.$ We use Lemma 2.19 to apply induction on the dimension of a scheme. See §2.1 below for more detailed outline of the proof.
@@
    Let $i:Z\to\mathfrak{X}$ be a closed immersion of codimension 11 and $Z$ an irreducible closed scheme of $\mathfrak{X}$ with $\operatorname{char}(Z)=p$. Then the canonical map induces a quasi-isomorphism
@@
For a scheme $X$, $X^{(s)}$ denotes the set of points of codimension $s$ and $k(X)$ denotes the ring of rational functions on $X$. For a ring AA, $k(A)$ denotes $k(\operatorname{Spec}(A))$. For a point $x\in X$, $\kappa(x)$ denotes the residue field of $x$, $\mathcal{O}_{X,x}$ denotes the local ring of $X$ at $x\in X$ and $\mathcal{O}^{h}_{X,x}$ denotes the henselization of $\mathcal{O}_{X,x}$.
@@ 2 Gersten-type conjecture
for some integer $a$ with $0\leq a\leq M=\operatorname{dim}(Y)$ (cf. [28, pp.180–181, Definition 2.1]). In this section, we prove the Gersten-type conjecture for a pair of a sheaf $\mathcal{F}$ on $Y_{\mathrm{\acute{e}t}}$ and a non-negative integer NN which satisfies Condition 2.11 over the henselization ring of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point $y\in Y$ ( Theorem 1.2, Theorem 2.22). Let
@@
The outline of the proof of Theorem 1.2 is as follows: By [17, p.93, III, Corollary 1.28], it suffices to prove $\operatorname{P}_{1}(a)$ and $\operatorname{P}_{2}(a)$ in the case where $Y$ is replaced by the spectrum of the henselization $A=\mathcal{O}^{h}_{Y,y}$ of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point $y\in Y$ and a point $y\in Y$ is replaced by the maximal ideal of AA. We put $\operatorname{Spec}(C)=\operatorname{Spec}(A)\times_{Y}Z$ and prove Theorem 1.2 by induction on $a=\#(\operatorname{Spec}(A)^{(0)})$.
@@ 2.2 Equi-characteristic cases
In this section, we frequently assume that a ring AA is a formally equidimensional noetherian local ring (cf. [16, p.251, Definition]). For example, an equidimensional local ring which is essentially of finite type over a regular local ring is formally equidimensional by [16, p.251, Theorem 31.6 (iii)].
@@
Let AA be a formally equidimensional noetherian local ring. By [16, p.251, Theorem 31.6 (iv)] and [16, p.250, Theorem 31.5], AA is equidimensional. By [16, p.251, Theorem 31.6 (i)], $A_{\mathfrak{p}}$ is also formally equidimensional for any $\mathfrak{p}\in\operatorname{Spec}(A)$. Moreover, the henselization $A_{\mathfrak{p}}^{h}$ of $A_{\mathfrak{p}}$ is formally equidimensional by the definition.
@@
Let AA be a formally equidimensional noetherian local ring, $\mathcal{F}$ a sheaf of groups on $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$ and $x\in\operatorname{Spec}(A)^{(s)}$. Put
@@
By Remark 2.3 and [17, p.93. III, Corollary 1.28], it suffices to prove the statement in the case where AA is a formally equidimensional henselian local ring and $x$ is the maximal ideal of $\operatorname{Spec}(A)$. We prove the statement by induction on $\operatorname{dim}(A)$.
@@
Let the notations AA, $\mathcal{F}$ and $T$ be the same as in Proposition 2.4. Suppose that $\operatorname{dim}(A)=s$. Then $\operatorname{H}^{t}_{\mathrm{\acute{e}t}}(A,\mathcal{F})$ is a torsion group for $t\geq s+T$.
@@
Let AA be a henselian local ring, II an ideal of AA and $\mathcal{F}$ a sheaf of groups on $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$. Then we have an isomorphism
@@
Let AA be a formally equidimentional noetherian local ring over a field $k$, $l$ a prime number which is prime to $\operatorname{char}(k)$ and $\mathcal{F}$ a sheaf of groups on $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$. Let the notation $T$ be the same as in Proposition 2.4. Suppose that $\operatorname{cd}_{l}(k)<\infty$ and $\operatorname{tr.deg}_{k}A_{x}<\infty$ for any $x\in\operatorname{Spec}(A)^{(0)}$. Put
@@
It suffices to prove the statement in the case where AA is a formally equidimensional henselian local ring and $x$ is the maximal ideal of $\operatorname{Spec}(A)$. We prove the statement by induction on $\operatorname{dim}(A)$.
@@
for $v\geq N+T+1$. In order to prove the statement, we use Lemma $2.$ Since AA is henselian, we have an isomorphism
@@
for $m\geq 0$ where $i:\operatorname{Spec}(\kappa(x))\to\operatorname{Spec}(A)$ is the closed immersion. Since AA is universally catenary by [16, p.251, Theorem 31.6 (iv)], we have
@@
Let AA be a formally equidimensional local ring of positive characteristic $p>0$ and $\mathcal{F}$ a sheaf of groups on $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$. Since $\operatorname{cd}_{p}(A)=1$, we have
@@
Let AA be a formally equidimensional local ring and $\mathcal{F}$ a sheaf on the étale site $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$. Let NN be a non-negative integer and $l$ a prime number. Suppose that $\operatorname{dim}(A)=s>1$ and $N\geq T$ where the notation $T$ is the same as in Proposition 2.4.
@@
for $s-1\leq e\leq s$. We prove the isomorphism (18) by using Lemma $2.$
@@
Let $\mathcal{V}$ be the category whose objects are normal crossing varieties over the spectrum of a field $k$ and whose morphisms are morphisms of schemes. Let $X$ be an object in $\mathcal{V}$, $\mathcal{F}_{X}$ an étale sheaf of abelian groups on $X$ and $f:X\to Y$ a morphism in $\mathcal{V}$. Let NN be a non-negative integer. Then $(\mathcal{F}_{X},N)\in\mathbb{S}_{X_{\mathrm{\acute{e}t}}}\times\mathbb{N}_{\geq 0}$ satisfies the followings:
@@
Let $X$ be essentially of finite type over the spectrum of a regular ring of dimension at most 11, $n$ an integer and $\mathbb{Z}(n)$ Bloch’s cycle complex for the Zariski topology. Let $i:Z\to X$ be a closed immersion of codimension $c$ with open complement $j:U\to X$, $i_{1}:W\to X$ a closed immersion of codimension $c^{\prime}$ and
@@
is exact for any integer ii. If $\operatorname{char}(k)=p\geq 0$, then we have a quasi-isomorphism
@@
for any integer ii. So the isomorphism (27) follows from Lemma 2.1. This completes the proof. ∎
@@
Let $Y$ be a smooth scheme over the spectrum of a field $k$ and $Z$ a normal crossing variety over $\operatorname{Spec}(k)$ which is a closed scheme of $Y$ of codimension 11.
Let AA be the henselization of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point $y$ of $Y$, $C$ the henselization of the local ring $\mathcal{O}_{Z,y}$ and $i:\operatorname{Spec}(C)\to\operatorname{Spec}(A)$ the corresponding closed immersion.
@@
We prove the injectivity of the homomorphism (37) for $t=N^{\prime}$ by induction on $s$. By [17, p.93, III, Corollary 1.28], it suffices to prove the injectivity of the homomorphism (37) in the case where AA is a henselian regular local ring with $\operatorname{dim}(A)=s$.
@@
In the case where $s=1$, we prove the injectivity of the homomorphism (37). Since AA is a henselian valuation ring, we have an isomorphism
@@
  • Property $\operatorname{P}_{3}(a)$: Let AA be the henselization of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point of $y\in Y$. Then the sequence
@@
First, we prove $\operatorname{P}_{1}(a+1)$. By [17, p.93, III, Corollary 1.28], it suffices to prove $\operatorname{P}_{1}(a+1)$ in the case where $Y$ is the spectrum of a henselian local ring AA and $\operatorname{dim}(Y)=s$. Since $(Z^{(0)})=a$, $\operatorname{P}_{1}(a+1)$ follows from $\operatorname{P}_{3}(a)$ and Lemma 2.18.
@@
is injective in the case where AA is the henselization at a point $y$ of $Y^{(s)}\cap Z$ by [17, p.93, III, Corollary 1.28]. Since we have
@@
Let $A_{a+1}$ be the henselization of $\mathcal{O}_{Y_{a+1},y}$ and $U^{\prime}=\operatorname{Spec}(A_{a+1})\times_{Y_{a+1}}U$. For notational simplicity, a base change of $j$ (resp. $j^{\prime}$, ii) is also denoted by $j$ (resp. $j^{\prime}$, ii), unless confusion arises. Then we have an isomorphism
@@
Finally, we prove $\operatorname{P}_{3}(a+1)$. Let the notations AA, $C$ and $U^{\prime}$ be the same as above. By (2.2) and $\operatorname{P}_{3}(a)$, the sequences
@@
are exact for $N^{\prime}\geq N_{1}$. Since AA is henselian, we have isomorphisms
@@
for $u+v=s^{\prime}+t$. Hence, the statement follows from Lemma $2.$ This completes the proof.
@@
Let AA be the strict henselian of a local ring of a normal crossing variety over the spectrum of a field of positive characteristic. Then AA is a direct limit of a family of the strict henselian rings $A_{i}$ ($i\in I$) of a local ring of a normal crossing variety over the spectrum of the prime field which satisfies the following:
Let $A/(a_{s})$ $(s\in S)$ be the irreducible components of AA. Then
@@
where $k_{i}$ run over subfields of $k$ which contain $k^{\prime}$ and are finitely generated over the prime field. Let $\mathfrak{m}$ be the maximal ideal of $C$, $f_{i}:k_{i}\otimes_{k^{\prime}}C^{\prime}\to C$ the base change and $\mathfrak{m}_{i}=(f_{i})^{-1}(\mathfrak{m}).$ Put $C_{i}=\left(k_{i}\otimes_{k^{\prime}}C^{\prime}\right)_{\mathfrak{m}_{i}}.$ Then $C_{i}$ is regular by [16, p.182, Theorem 23.7 (i)]. Let $c^{\prime}_{s}$ ($s\in S$) be elements of $C$ such that $C/(c^{\prime}_{s})\simeq A^{\prime}/(a^{\prime}_{s})$ where $A^{\prime}/(a^{\prime}_{s})$ ($s\in S$) are the irreducible components of $A^{\prime}$. Since we are able to choose a directed subset $J$ of II such that $c_{s}^{\prime}\in\mathfrak{m}_{j}$ and
@@
for any $i\in I$ and AA (resp. $A_{i}$) denotes the strict henselian of $A^{\prime}$ (resp. $A^{\prime}_{i}$). Then AA is also the strict henselian of $\mathcal{O}_{Y,y}$ and $A_{i}$ satisfies the property of the statement. This completes the proof. ∎
@@
Let AA be the henselization of the local ring $\mathcal{O}_{Y,y}$ of a normal crossing variety $Y$ over the spectrum of a field of positive characteristic $p>0$ at a point $y\in Y$. Then the sequence